In a bakery, 3 cakes of fresh cream pineapple, 4 cakes of chocolate, 2 cakes of buttercream strawberryand 1 cake without cream are available. If two customers purchase one cake each, such that first cakeis replaced before the sale of the second then, the probability that both the cakes sold would be ofchocolate flavor is?​

1
Expert's answer
2021-03-30T08:07:08-0400

Number of possible outcomes = 3+4+2+1 = 10

P(Chocolate cake) = 4/10 = 0.4

Let the events:

A = first customer purchases a chocolate cake

B = second customer purchases a chocolate cake

Since the two events are independent,

P(A and B) = P(A) * P(B) = 0.4*0.4 = 0.16


Answer = 0.16